OMEGA(1) OMEGA(1) NAME omega,iniomega,viromega−extendedunicodeTeX SYNOPSIS omega[options][commands] DESCRIPTION Thismanualpageisnotmeanttobeexhaustive.Thecompletedocumentationforthisversionof TEXcanbefoundintheinfofileormanualWeb2C:ATeXimplementation. OmegaisaversionoftheTEXprogrammodifiedformultilingualtypesetting.Itusesunicode, andhasadditionalprimitivesfor(amongotherthings)bidirectionaltypesetting. TheiniomegaandviromegacommandsareOmega’sanaloguestotheinitexandvirtexcommands.Inthisinstallation,theyaresymlinkstotheomegaexecutable. Omega’scommandlineoptionsaresimilartothoseofTEX. Omegaisexperimentalsoftware. OPTIONS ThisversionofOmegaunderstandsthefollowingcommandlineoptions. --fmtformat Useformatasthenameoftheformattobeused,insteadofthenamebywhichOmega wascalledora%&line. --helpPrinthelpmessageandexit. --iniBeiniomega,fordumpingformats;thisisimplicitlytrueiftheprogramiscalledas iniomega. --interactionmode Setstheinteractionmode.Themodecanbeoneofbatchmode,nonstopmode,scrollmode, anderrorstopmode.Themeaningofthesemodesisthesameasthatofthecorresponding \commands. --ipcSendDVIoutputtoasocketaswellastheusualoutputfile.Whetherthisoptionis availableisthechoiceoftheinstaller. --ipc-start As--ipc,andstartstheserverattheotherendaswell.Whetherthisoptionisavailable isthechoiceoftheinstaller. --kpathsea-debugbitmask Setspathsearchingdebuggingflagsaccordingtothebitmask.SeetheKpathseamanual fordetails. --maketexfmt Enablemktexfmt,wherefmtmustbeoneoftexortfm. --no-maketexfmt Disablemktexfmt,wherefmtmustbeoneoftexortfm. --output-commentstring UsestringfortheDVIfilecommentinsteadofthedate. --prognamename Pretendtobeprogramname.Thisaffectsboththeformatusedandthesearchpaths. --shell-escape Enablethe\write18{command}construct.ThecommandcanbeanyBourneshellcommand.Thisconstructisnormallydisallowedforsecurityreasons. --version Printversioninformationandexit. Web2C7.32x 27December1997 1 OMEGA(1) OMEGA(1) ENVIRONMENT SeetheKpathsearchlibrarydocumentation(the‘Pathspecifications’node)forprecisedetailsof howtheenvironmentvariablesareused.Thekpsewhichutilitycanbeusedtoquerythevalues ofthevariables. Onecaveat:InmostOmegaformats,youcannotuse˜ inafilenameyougivedirectlytoOmega, because˜ isanactivecharacter,andhenceisexpanded,nottakenaspartofthefilename.Other programs,suchasMETAFONT,donothavethisproblem. TEXMFOUTPUT Normally,Omegaputsitsoutputfilesinthecurrentdirectory.Ifanyoutputfilecannot beopenedthere,ittriestoopenitinthedirectoryspecifiedintheenvironmentvariable TEXMFOUTPUT.Thereisnodefaultvalueforthatvariable.Forexample,ifyousay texpaperandthecurrentdirectoryisnotwritable,ifTEXMFOUTPUThasthevalue /tmp,Omegaattemptstocreate/tmp/paper.log(and/tmp/paper.dvi,ifanyoutputis produced.) TEXINPUTS Searchpathfor\inputand\openinfiles.Thisshouldprobablystartwith‘‘.’’,sothat userfilesarefoundbeforesystemfiles.Anemptypathcomponentwillbereplacedwith thepathsdefinedinthetexmf.cnffile.Forexample,setTEXINPUTSto ".:/home/usr/tex:"toprependthecurrentdirecoryand‘‘/home/user/tex’’tothestandardsearchpath. TEXEDIT Commandtemplateforswitchingtoeditor.Thedefault,usuallyvi,issetwhenOmegais compiled. FILES Thelocationofthefilesmentionedbelowvariesfromsystemtosystem.Usethekpsewhichutilitytofindtheirlocations. omega.pool EncodedtextofOmega’smessages. *.fmtPredigestedOmegaformat(.fmt)files. BUGS ThisversionofOmegafailstotraparithmeticoverflowwhendimensionsareaddedorsubtracted. Caseswherethisoccursarerare,butwhenitdoesthegeneratedDVIfilewillbeinvalid. TheDVIfilesproducedbyOmegamayuseextensionswhichmakethemincompatiblewithmost softwaredesignedtohandleDVIfiles.Inordertoprintorpreviewthem,youshoulduseodvips togenerateaPostScriptfile. Omegaisexperimentalsoftware.Ifyouuseit,subscribetotheomegamailinglistomega@ens.fr bysendingamessagecontainingsubscribeomegaYourNametolistserv@ens.fr. SEEALSO tex(1),mf(1),odvips(1), AUTHORS TheprimaryauthorsofOmegaareJohnPlaiceandYannisHaralambous. Web2C7.32x 27December1997 2